# Do LLMs Have Distinct and Consistent Personality? TRAIT: Personality Testset designed for LLMs with Psychometrics

### v1

@@ Abstract
The idea of personality in descriptive psychology, traditionally defined through observable behavior, has now been extended to Large Language Models (LLMs) to better understand their behavior. This raises a question: do LLMs exhibit distinct and consistent personality traits, similar to humans? Existing self-assessment personality tests, while applicable, lack the necessary validity and reliability for precise personality measurements.
To address this, we introduce TRAIT, a new tool consisting of 8K multi-choice questions designed to assess the personality of LLMs with validity and reliability. TRAIT is built on the psychometrically validated human questionnaire, Big Five Inventory (BFI) and Short Dark Triad (SD-3), enhanced with the ATOMIC10$\times$ knowledge graph for testing personality in a variety of real scenarios. TRAIT overcomes the reliability and validity issues when measuring personality of LLM with self-assessment, showing the highest scores across three metrics: refusal rate, prompt sensitivity, and option order sensitivity. It reveals notable insights into personality of LLM: 1) LLMs exhibit distinct and consistent personality, which is highly influenced by their training data (i.e., data used for alignment tuning), and 2) current prompting techniques have limited effectiveness in eliciting certain traits, such as high psychopathy or low conscientiousness, suggesting the need for further research in this direction [^1] The code is available at github.com/pull-ups/TRAIT.
**footnotetext: equal contribution
(a) Comparing self-assessment tests and our TRAIT.
(b) TRAIT has high validity, reliability, and human qualification results.
Figure 1: TRAIT is a personality test for LLMs based on trusted questionnaires John et al. (1999); Jones and Paulhus (2014) and large-scale commonsense knowledge graphs to cover wide range of real-world scenarios West et al. (2022). TRAIT provides higher reliability (e.g., low prompt sensitivity) and validity (low refusal rate) (§2), and achieves 98.0% accuracy when validated by human experts (§3).
@@ 1 Introduction
“We are what we repeatedly do.”
Durant, 1927
Descriptive psychology defines personality as observable fact measuring behavior Bergner (2020); Jeffrey (1990); Putman (1990). Just as we consider someone assertive who often speaks in a commanding tone, researchers in psychology have measured and scored one’s personality as an enduring pattern of behavior and linguistic output, not as an inner mechanism nor a causal entity.
Figure 2: (Left) Responses from LLMs to self-assessment tests (e.g., BFI) fail to capture the personality of the models in various real scenarios pertaining to personality. (Right) TRAIT covers a wide range of contexts to better portray the personality traits of LLMs.
@@
As Large Language Models (LLMs) become increasingly intelligent and more closely integrated into human life, the concept of personality has been extended to these models to better understand their behavioral patterns. Do LLMs exhibit distinct and consistent behavioral patterns for various contexts and inputs, similar to humans?
To address these questions, we present TRAIT (TRait of AI Testbench), a reliable and valid questionnaire designed to assess personality traits of LLMs. Our work aims to shed new light on patterning the responses of LLMs and further suggest potential approaches for employing LLMs in many real-world applications Ammanabrolu et al. (2022).
However, it is challenging to accurately measure the personality of language models. As illustrated in Figure 1(a), commonly used self-assessment personality tests, such as Big Five Inventory (BFI) John et al. (1991), Anthropic Personality Eval (Anthropic-Eval) Perez et al. (2022)), ask for LLMs to introspect and report itself about the statement. As these assessments only focus on responses to general questions (e.g., “Are you talkative?”) rather than context-specific ones (e.g., “Are you talkative when greeting new friends?”), this approach may not accurately capture how LLMs behave in actual situations. Also, as in Figure 1(b), the results of self-assessment sway depending on the prompt format and show a high rate of refusal, which compromises the reliability and validity of measurement (§2).
@@
Based on these findings, we introduce TRAIT, the first LLM personality test which considers diverse aspects of reliability and validity in psychometrics, to the best of our knowledge. For the data construction, we collect 71 validated questionnaire items from human assessments — BFI and Short Dark Triad (SD-3) Jones and Paulhus (2014) — as our seed dataset. We further enrich them to unique detailed scenarios with ATOMIC10$\times$ West et al. (2022), a large-scale commonsense knowledge graph. TRAIT includes 8,000 items, which is 112 times larger compared to the seed dataset which enables us to draw statistically significant conclusions about the LLMs’ responses and behavior patterns in various realistic contexts (§3).
@@
In our analysis of nine LLMs using TRAIT, we make three key observations related to the personality of LLMs (§4): 1) LLMs display statistically distinctive and consistent behavioral patterns. For instance, GPT-4 is significantly more agreeable than GPT-3.5. 2) Alignment tuning [^2] Alignment tuning here is an overarching term for SFT, RLHF, and RLAIF Lin et al. (2023). alters the LLMs’ personality across various traits: it decreases extraversion, openness, and socially adversarial traits (Dark Triad), and increases agreeableness and conscientiousness. 3) Prompting can induce specific personality in LLM, however, it can not elicit certain traits, e.g., high level of psychopathy. We will publicly release our TRAIT to establish a foundation for understanding the personality of LLMs and to guide these models to align their behavior with human values.
@@ 2 Measuring Personality of LLM
Here, we review how previous works measure LLMs’ personality, and empirically show that self-assessment personality tests lack acceptable validity and reliability when measuring the personality of LLMs. These findings motivated us to develop TRAIT, a personality test designed for LLMs with high validity and reliability.
@@ 2.1 Big Five and Dark Triad
Dataset Validity Reliability
Refusal Rate ($\downarrow$) Prompt Sens. ($\downarrow$) Option Order Sens. ($\downarrow$) Paraphrase Sens. ($\downarrow$) Avg. Sens. ($\downarrow$)
BIG-5 38.2 $\pm$ 1.32 40.9 $\pm$ 5.29 65.5 $\pm$ 2.53 18.8 $\pm$ 2.97 41.7
SD-3 41.8 $\pm$ 2.60 48.1 $\pm$ 6.07 62.1 $\pm$ 3.61 20.3 $\pm$ 4.63 43.5
IPIP-NEO-PI 39.6 $\pm$ 0.66 41.5 $\pm$ 1.83 63.1 $\pm$ 1.01 20.4 $\pm$ 1.36 41.7
Anthropic-Eval 19.8 $\pm$ 0.15 30.1 $\pm$ 0.51 40.5 $\pm$ 0.32 32.6 $\pm$ 0.64 34.4
TRAIT (Ours) 0.2 $\pm$ 0.01 26.0 $\pm$ 0.51 29.3 $\pm$ 0.35 20.1 $\pm$ 0.38 25.1
Table 1: Validity and reliability of LLM personality tests. For each cell, we average the metric from 7 different models, jointly with the 95% confidence interval of the standard deviation. See Table 131516 for all results.
There are various frameworks to analyze the complex concept of personality. In our study, we adopt the most widely utilized frameworks for human personality analysis in the psychology literature; Dark Triad Paulhus (2014) and Big Five (BIG-5) McCrae and Costa Jr (1987); Gosling et al. (2003). Dark Triad comprises three socially adverse traits: Machiavellianism, Narcissism, and Psychopathy. BIG-5 identifies personality dimensions with five traits: Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism. Table 2 includes eight traits and facets covered in this paper. See Appendix C.2 for more details on these frameworks.
@@ 2.2 Existing Self-assessment Personality Tests
We assess four personality tests that are used on LLMs in previous studies. Three are well-established self-assessment333 Self-assessment, where individuals evaluate their personality, is commonly used in measuring human personality due to its simplicity. Alternatively, personality can be inferred from observing patterns in behavior, a method called behavioral and performance measures, or objective personality testing Ortner and Proyer (2015). More related works are in Appendix C.2. tests which are designed to measure personality of humans: BFI John et al. (1991) (44 items), SD-3 Jones and Paulhus (2014) (27 items) and IPIP-NEO-PI Goldberg et al. (1999) (300 items). These tests are recognized for their reliability and validity when testing human personality as they are crafted by psychology experts, and these are often used to measure LLMs’ personality as well Serapio-García et al. (2023). However, the number of items in the tests is limited, and the effectiveness of these tests for LLMs is questionable since the answer to the self-assessment may not assert an LLM’s behavior in real-world scenarios. Additionally, we examine Anthropic-Eval Perez et al. (2022), a LLM-generated test specifically developed for evaluating LLMs’ personality. This test is also a self-assessment test, featuring 8,000 yes/no questions each accompanied by a label that reflects the response consistent with the assessed personality. See Table 3 for more statistics about the tests.
2.3 Validity and Reliability of Self-assessment Personality Tests
Validity and reliability are key ideas in psychometrics for confirming the quality of tests. Validity refers to how well the test measures what it is intended to measure. Reliability measures how the instrument produces similar results in different conditions Roberts and Priest (2006).
Validity metric.
We evaluate the validity by refusal rate, which calculates the ratio of how LLM refuses the given queries. A high refusal rate can obstruct the fair comparison among the individual models, potentially distorting the intended measurement and reducing its validity.
@@ 2.3 Validity and Reliability of Self-assessment Personality Tests
We assess reliability with three metrics which are motivated by test-retest reliability and parallel-form reliability in psychometrics. Test-retest reliability explains the agreement between the results of successive measurements of the same measure. For test-retest reliability, we define a prompt sensitivity, by adopting three different prompt templates from prior works Jiang et al. (2024); Miotto et al. (2022); Huang et al. (2023), and assessing if the three responses on each test items are not same. Additionally, we introduce option-order sensitivity by changing the order of options, and assessing if changing the order of options affects the response. For Likert-type QA, we reverse the option order starting with “very disagree” from “very agree” in the original form.
Parallel-form reliability explains a correlation derived from responses to two different versions of the test APA (2018). Inspired by parallel-form reliability in psychometrics, we measure paraphrase sensitivity, where we see if the same answer is given to a question with the same meaning but lexically different, by counting the mismatched answers between the original and paraphrased queries and calculate their ratio. See more details on these metrics in Appendix E.1.
@@
We assess validity and reliability of existing personality for LLMs, using seven different models for the assessment444 GPT-4, GPT-3.5, Mistral-7B-instruct, Mistral-7B-sft, Llama3-8B-instruct, Llama2-7B-chat, Tulu2-7B-DPO.. The results are shown in Table 1, with two findings: 1) Personality tests for humans have a surprisingly high refusal rate and low reliability when testing LLM personality. BIG-5, SD-3, and IPIP-NEO-PI all show a high refusal rate, indicating that at least 38.2%, LLMs refuse to answer when asked to assess their personality. It also shows low reliability when measured by three different sensitivities: on average, all these tests show 41.7%-43.5% of sensitivity, which means the personality scores of models can easily fluctuate with minor changes in the format of the tests. 2) Anthropic-Eval, a personality test designed for LLMs, has better validity and reliability than previous three tests, but yet not adequate. It shows a refusal rate of 19.8% and an average sensitivity of 34.4%, which are both improved numbers when compared to all personality tests for humans. However, it still has a non-marginal refusal rate and as shown from the highest paraphrase sensitivity of 32.6%, reliability also has non-trivial room for improvement.
@@ 3 TRAIT: Reliable and Valid LLM Personality Tests
Dataset #Traits #Items Dist-3 ($\uparrow$) Ent-3 ($\uparrow$) Avg. Question Len. Assessment Detailed Scenario
SD-3 Jones and Paulhus (2014) 3 27 - - 28.8 Likert
BFI John et al. (1999) 5 44 - - 46.0 Likert
IPIP-NEO-PI Goldberg et al. (1999) 5 300 - - 26.8 Likert
Anthropic-Eval Perez et al. (2022) 8 8,000 0.529 14.1 52.2 Yes/No
TRAIT (Ours) 8 8,000 0.618 17.4 766.7 Multi-choice
Table 3: Dataset statistics. Dist-3 and Ent-3 are the metrics for the lexical diversity Han et al. (2022a), and we do not include numbers from human assessments (SD-3, BFI, IPIP-NEO-PI-300) due to their small size. See Table 7 for representative examples.
We thus develop TRAIT, a new multi-dimensional personality test to assess LLM’s personality on eight traits from Dark Triad and BIG-5. For better validity and reliability, TRAIT includes: 1) more comprehensive semantic diversity — expanded from 71 small, validated human self-assessments to 112 times larger dataset (§3.1), and 2) detailed guideline to allow any model available for multi-choice question-answering (§3.2).
@@ 3.1 Dataset Construction Pipeline
We construct TRAIT with Human-AI collaboration. All the prompts used to condition GPT-4 when constructing data are in Appendix K.
@@
We start by collecting 71 items of self-assessment questionnaires: 44 items from BFI and 27 from SD-3. To create detailed and varied descriptions of personality, we use GPT-4 to generate 240 unique descriptions based on 8 to 10 collected questionnaires. We then filter out 40 sentences that either have high Jaccard similarity with others or are deemed inaccurate, resulting in 1,600 diverse sentences.
@@
We show in §2.3 that self-assessment tests have a high refusal rate and low reliability, and we suspect the general question that asks for self-reporting (e.g., “Are you talkative”) in self-assessment forms is the main reason. So, we expand these 1,600 personality descriptions into 8,000 more detailed scenarios that can have personality-induced decisions as plausible action space. We use ATOMIC10$\times$ West et al. (2021), a large commonsense knowledge graph with 6.45 million entries, including a wide range of physical and social situations (e.g., if X and Y argue, so, X wants to (xWant) avoid Y). Given each personality description, we randomly sample 20 situations from ATOMIC10$\times$, and then pick the five most relevant ones using GPT-4. Concurrently, we induce GPT-4 to craft a situation and question given the personality description and situation from ATOMIC10$\times$.
@@
Finally, for each detailed scenario, we create a multiple-choice question with four options. Two of these options are likely to be selected by respondents with a strong presence of the trait (High), while the other two are more likely to be chosen by those with a weaker presence of the trait (Low). This helps us to various potential responses to the scenarios, covering a balanced facet of each personality trait (see Appendix D.2.1 for more details).
3.2 Measuring Personality Scores with TRAIT using Token Probability
We follow the evaluation protocol of existing multi-choice question-answering (MCQA) benchmarks such as MMLU Hendrycks et al. (2020) which uses token probabilities of the four options for evaluation. To mitigate bias from the order of the options, we alternate the arrangement of options twice and averaged. More details are in Appendix D.2.

### v2

@@ Abstract
Recent advancements in Large Language Models (LLMs) have led to their adaptation in various domains as conversational agents. We wonder: can personality tests be applied to these agents to analyze their behavior, similar to humans? We introduce TRAIT, a new benchmark consisting of 8K multi-choice questions designed to assess the personality of LLMs. TRAIT is built on two psychometrically validated small human questionnaires, Big Five Inventory (BFI) and Short Dark Triad (SD-3), enhanced with the ATOMIC10$\times$ knowledge graph to a variety of real-world scenarios. TRAIT also outperforms existing personality tests for LLMs in terms of reliability and validity, achieving the highest scores across four key metrics: Content Validity, Internal Validity, Refusal Rate, and Reliability. Using TRAIT, we reveal two notable insights into personalities of LLMs: 1) LLMs exhibit distinct and consistent personality, which is highly influenced by their training data (e.g., data used for alignment tuning), and 2) current prompting techniques have limited effectiveness in eliciting certain traits, such as high psychopathy or low conscientiousness, suggesting the need for further research in this direction.111 [Uncaptioned image] Code: pull-ups/TRAIT [Uncaptioned image] Data: mirlab/TRAIT
Figure 1: TRAIT is a personality test for LLMs based on trusted questionnaires John et al. (1999); Jones and Paulhus (2014) and large-scale commonsense knowledge graphs West et al. (2022). LLMs show discrepancy in self-assessing their personality and actual decision making.
@@ 1 Introduction
Just as we consider someone assertive who often speaks in a commanding tone, researchers in psychology have measured one’s personality as an enduring pattern of behavior and linguistic output, not as an inner mechanism nor a causal entity Bergner (2020). As Large Language Models (LLMs) become more closely integrated into human life, the concept of personality can be extended to better understand their behavioral patterns Perez et al. (2022). Do LLMs exhibit distinct and consistent behavioral patterns for various contexts and inputs, similar to humans?
Answering this question requires a reliable and valid test set to measure the personality of LLMs. However, existing questionnaires ask subjects to introspect and report about the statement Li et al. (2024); Wang et al. (2024), i.e., self-assessment tests, which lacks reliability and validity. As illustrated in Figure 1, such general questions (e.g., “Are you outgoing and sociable?”) may not accurately capture how LLMs behave in actual situations (e.g., “What hobby would you recommend?”). Moreover, LLM responses vary significantly with details of prompting (Sclar et al., 2024) and often include refusals, undermining the reliability and validity of measurement (§2).
@@
Based on these findings, we present TRAIT (TRait of AI Testbench), a reliable and valid questionnaire designed to assess personality traits of LLMs. Our work aims to shed new light on patterning the responses of LLMs and further suggest potential approaches for employing LLMs in many real-world applications Ammanabrolu et al. (2022). For the data construction, question items from the widely recognized questionnaire, BFI John et al. (1999) and Short Dark Triad (SD-3) Jones and Paulhus (2014), were used as seeds for augmentation. Then we further enrich them to unique detailed scenarios with ATOMIC10$\times$ West et al. (2022), a large-scale commonsense knowledge graph. TRAIT includes 8,000 items, which is 112 times larger compared to the seed dataset which enables us to draw statistically significant conclusions about the LLMs’ responses and behavior patterns in various realistic contexts (§3).
@@
In our analysis of nine state-of-the-art LLMs using TRAIT, we make three key observations related to the personality of LLMs (§4): 1) LLMs display statistically distinctive and consistent behavioral patterns. For instance, GPT-4 is significantly more agreeable than GPT-3.5. 2) Alignment tuning [^2] Alignment tuning here is an overarching term for SFT, RLHF, and RLAIF Lin et al. (2023). alters the LLMs’ personality across various traits: it decreases extraversion, openness, and socially adversarial traits (Dark Triad), and increases agreeableness and conscientiousness. 3) Prompting can induce specific personality in LLM, however, it can not elicit certain traits, e.g., high level of psychopathy. We will publicly release our TRAIT to establish a foundation for understanding the personality of LLMs and to guide these models to align their behavior with human values.
@@ 2 Measuring Personality of LLM
Here, we review how previous works measure LLMs’ personality, and empirically show that self-assessment personality tests lack reliability and validity when measuring the personality of LLMs. These findings motivate us to develop TRAIT, a personality test designed for LLMs with high reliability and validity.
@@ 2.1 Big Five and Dark Triad
Dataset #Items Dist-3 ($\uparrow$) Assessment Detailed Scenario
SD3 27 - Likert
BFI 44 - Likert
IPIP-NEO-PI 300 - Likert
Anthropic-Eval 8,000 0.529 Likert
Our Dataset 8,000 0.618 Multi-choice
Table 2: Dataset statistics. Dist-3 is a metric for lexical diversity. See Table 8 for all representative examples of SD3 Jones and Paulhus (2014), BFI John et al. (1999), IPIP-NEO-PI Goldberg et al. (1999), and Anthropic-Eval Perez et al. (2022).
Content Val. ($\uparrow$) Internal Val. ($\uparrow$) Refusal ($\downarrow$) Reliability ($\downarrow$)
Diversity. Score Diff. Generation MCQ Prompt Option Order Paraphrase Avg.
BIG-5 BFI* - 45.0 53.9 30.8 37.2 62.0 22.9 40.7
IPIP-NEO-PI* - 40.0 49.5 28.1 44.5 62.3 24.5 43.8
Anthropic-Eval* 61.1 62.5 41.7 17.4 27.2 36.7 27.1 30.3
TRAIT 71.9 77.5 3.1 0.0 31.6 33.5 24.5 29.8
Dark Triad SD-3* - 33.3 45.7 27.7 54.7 66.5 27.3 49.5
Anthropic-Eval* 45.3 41.6 40.6 14.8 33.9 40.2 32.4 35.5
TRAIT 51.0 83.3 3.3 0.0 28.1 28.2 16.8 24.4
*denotes questionnaire based on Likert scale assessment, and denotes questionnaire based on multi-choice question assessment.
Table 3: Validity score, Refusal rate and reliability score of LLM personality tests. Each cell shows the average metric from 8 different models. TRAIT demonstrates the lowest refusal rate while showing the highest validity and average of reliability. Generation and MCQ in Refusal indicate refusal on open-generation and multiple-choice question setting respectively. Prompt, Option Order, and Paraphrase in Reliability indicate sensitivity on prompt, option order and paraphrase, respectively. See Table 14 for confidence interval and Table 151718 for all results.
There are various frameworks to analyze the complex concept of personality. In our study, we adopt the most widely utilized frameworks for human personality analysis in the psychology literature; Dark Triad Paulhus (2014) and Big Five (BIG-5) McCrae and Costa Jr (1987); Gosling et al. (2003). Dark Triad comprises three socially adverse traits: Machiavellianism, Narcissism, and Psychopathy. BIG-5 identifies personality dimensions with five traits: Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism. Table 1 includes eight traits and their facets. See Appendix C.2 for more details on these frameworks.
@@ 2.2 Existing Self-assessment Personality Tests
We assess four personality tests that are used on LLMs in previous studies. Three are well-established self-assessment333 Self-assessment, where individuals evaluate their personality, is commonly used by human subjects due to its simplicity. Alternatively, there is another method called ‘behavioral and performance measures’, or ‘objective personality testing’ Ortner and Proyer (2015) which infers personality from observing patterns in behavior. More related works are in Appendix C.2. tests which are designed to measure personality of humans: BFI John et al. (1991) (44 items), SD-3 Jones and Paulhus (2014) (27 items) and IPIP-NEO-PI Goldberg et al. (1999) (300 items). These tests are recognized for their reliability and validity when testing human personality as they are crafted by psychology experts, and these are often used to measure LLMs’ personality as well Serapio-García et al. (2023). However, the number of questions is limited, ranging from 27 to 300, and the effectiveness of these tests for LLMs is questionable since the answer to the self-assessment may not assert an LLM’s behavior in real-world scenarios. Additionally, we examine Anthropic-Eval Perez et al. (2022), a LLM-generated test specifically developed for evaluating LLMs’ personality. This test is also a self-assessment test, featuring 8,000 binary (yes/no) questions. See Table 2 for more statistics about the tests.
2.3 Assessing the Quality of Personality Tests
The appropriateness and robustness of test questionnaires are often evaluated with Validity and reliability in human psychometrics Roberts and Priest (2006). We additionally report response refusal rate to analyze LLM-specific test failure cases.
Validity metrics.
Content validity assesses how comprehensively a test measures its intended concept. As a good personality questionnaire should include all the facets evenly, we annotate each item in the questionnaire to indicate its associated facet (or the facet it aims to measure) from the list in Table 1 using GPT-4  [^4] We used gpt-4o-2024-08-06 for evaluations.. Based on the annotations, we measure the 3 diversity indices: Simpson Diversity ndex Simpson (1949), Shannon-Wiener index Shannon (1948), and Evenness Index Pielou (1966) (Details in Appendix E.1). TRAIT achieves higher scores than Anthropic-Eval across all personality traits, which indicates a more even distribution of facets in the questionnaire, with both being LLM-generated on the same scale.
On personality assessment, internal validity refers to the causal relationship between the independent variable and the dependent variable. While keeping all other conditions the same, we compared test scores obtained after assigning personalities at different levels through prompting (e.g., “You are an assistant with high/low extraversion.”). Then we reported the proportion of cases where increase or decrease was applied in the test score according to the induced personality.
@@ 2.3 Assessing the Quality of Personality Tests
Paraphrase sensitivity is inspired by the concept of parallel-form reliability in psychometrics. Parallel-form reliability measures how consistent results are when using two different versions of the same test APA (2018). We use GPT-3.5 to create twin questions - paraphrases that have the same meaning as the original questions but use different words. Then we count how many times the answers differ between the original and paraphrased questions as a ratio. This helps us understand how sensitive the model’s responses are to changes in wording, even when the meaning stays the same. See more details in Appendix E.1.
Additionally, prompt sensitivity and option order sensitivity are representative types of sensitivity displayed by LLMs in MCQ settings. For prompt sensitivity, with a similar format with prior works Jiang et al. (2024); Miotto et al. (2022); Huang et al. (2023), we measure the consistency of answering across three instruction prompt templates as a ratio. For option order sensitivity, we evaluated whether swapping the position of high option and low option affects the consistency of responses. For Likert-type QA, we reversed the option order, starting with “very disagree” instead of “very agree” in the original form.
Refusal Rate.
We assess the refusal rate - the frequency of query rejections. High refusal rates can hinder fair model comparisons, potentially compromising measurement validity and reliability. We evaluate refusal rate on both open-ended and multiple-choice formats, and the criteria for determining whether a response is a refusal or not are presented in Appendix G.1.
@@
We assess refusal rate and reliability of LLMs responses, using eight different models555 GPT-3.5, Mistral-7B-instruct, Mistral-7B-sft, Llama3-8B-instruct, Tulu2-7B, Tulu2-7B-DPO, Gemma-2B-it, OLMo-7B-sft. The results are shown in Table 3, with two findings:
1) Personality tests for humans have a surprisingly high refusal rate when testing with LLMs. LLMs refused to answer nearly half of the questions on self-assessment tests (BIG-5, SD-3, and IPIP-NEO-PI). We found that this phenomenon did not significantly improve even when attempting to circumvent direct answers in the form of MCQ. We speculate that the introspective and self-reporting nature of Human Questionnaires is the direct cause of the high refusal rate Li et al. (2024); Wang et al. (2024).
2) Personality measurement based on Likert scales using human questionnaires has shown low scores in various aspects of robustness. The Likert scale-based measurement requires an additional process: projecting the respondent’s inner thoughts about a given context onto an appropriate scale. We suspect that language models particularly struggle with projecting scores robustly, as it is known that they are highly sensitive to slight changes in context Gupta et al. (2024).
@@ 3 TRAIT: Reliable and Valid LLM Personality Tests
We thus develop TRAIT, a new multi-dimensional personality test to assess LLM’s personality on eight traits from Dark Triad and BIG-5. For better validity and reliability, TRAIT includes: 1) more comprehensive semantic diversity — expanded from 71 small, validated human questionnaire items to 112 times larger dataset (§3.1), and 2) detailed guideline to allow any model available for multi-choice question-answering (§3.4).
@@ 3.1 Dataset Construction Pipeline
All the prompts used to condition GPT-4 when constructing data are in Appendix. As shown in Figure 2, we construct TRAIT with Human-AI collaboration. Detailed example is in Table 4, and prompts we use are in Appendix  K.
Processing Stage Example (Ext)
Self-assessment I’m outgoing and sociable.
Diverse Personality Description I prefer sociable hobbies to quiet, solitary ones.
Detailed Scenarios I walk to clear my mind. Inviting friends makes it social instead of peaceful. What should I do?
Multi-choice A. Start solo walks.
B. Maintain social walks.
C. Start a new quiet, solitary hobby.
D. Invite a new friend to join a mindfulness class together.
Table 4: Example of Dataset Making Process. Example sentences are condensed due to page limitations.
@@
The human questionnaire items, while capturing core components of personality, are too brief and summarized (e.g., I am talkative.) to cover diverse aspects of personality. For example, “Talkative” can manifest in diverse aspects, such as quantity of speech or initiation of conversations. To address this, we use GPT-4 to expand the 71 BFI and SD-3 items into 1,600 diverse personality descriptions across 8 personality types.
@@
While the 1,600 personality descriptions depict characteristics of traits expansively, we aim to simulate real-life decision-making more closely by developing specific contexts where subjects interpret situations and make judgments. We augment our dataset to 8,000 context-rich user queries. We use ATOMIC10$\times$ West et al. (2022), a large commonsense knowledge graph with 6.45 million entries, including a wide range of physical and social situations (e.g., X and Y argue, so, X wants to (xWant) avoid Y). Given each personality description, we randomly sample 20 situations from ATOMIC10$\times$, and then pick the five most relevant ones using GPT-4. Concurrently, we induce GPT-4 to craft a situation and question given the personality description and situation from ATOMIC10$\times$.
@@
Finally, for each detailed scenario, we create a multiple-choice question with four options. Two of these options are likely to be selected by respondents with a strong presence of the trait (High), while the other two are more likely to be chosen by those with a weaker presence of the trait (Low). This helps us to embrace various potential responses to the scenarios, covering a balanced facet of each personality trait (see ‘Content Validity’ in Table 3).